Makana invested $700 in a simple interest account that had an interest rate that was 3% more than that of her friend Marlys. If Marlys earned $39 after one year from an investment of $650, how much did Makana earn in one year?
Answer by richwmiller(17219)

700*(x+3)/100=y
7x+21=y
650*x/100=39
6.5x=39
x=39/6.5=6 marlys rate of interest
7x+21=y
42+21=y
y=63 Makana earned